OpenStudy (breanna.rae):

Algebra 1B. How many solution does the system have? 2x=-10y+6 and x+5y=3

OpenStudy (vrajrpatel):

x+5y=3;2x=−10y+6 Step: Solvex+5y=3for x: x+5y+−5y=3+−5y(Add -5y to both sides) x=−5y+3 Step: Substitute−5y+3forxin2x=−10y+6: 2x=−10y+6 2(−5y+3)=−10y+6 −10y+6=−10y+6(Simplify both sides of the equation) −10y+6+10y=−10y+6+10y(Add 10y to both sides) 6=6 6+−6=6+−6(Add -6 to both sides) 0=0

OpenStudy (breanna.rae):

Thank you!

OpenStudy (breanna.rae):

can you help me with another? @vrajrpatel